What is the theoretical yield of sodium chloride formed from the reaction of 2.2g of hydrochloric acid and 3.4g of sodium hydroxide?



1
Expert's answer
2021-11-13T01:36:39-0500

2.2+3.4=5.6

= 2.2/5.6×100= 39.29%